OpenStudy (chycora):

What is the area of the right triangle above? 144 in2 18 in2 72 in2 36 in2

OpenStudy (anonymous):

Pic

OpenStudy (anonymous):

Is there a picture to go along with the question?

OpenStudy (anonymous):

What are the dimension just tell me

OpenStudy (chycora):

the hight is 18 and the lengh is 4

OpenStudy (anonymous):

36

OpenStudy (anonymous):

36in2

OpenStudy (anonymous):

fan and medal and @ for more

OpenStudy (anonymous):

\[Area~of~\triangle=\frac{ 1 }{ 2 }\times~length~\times~height\]